An archaeologist discovers a wooden mask whose carbon-14 to carbon-12 ratio is one- sixteenth the ratio measured in a newly fallen tree. How old does the wooden mask seem to be, given this evidence and the ½-life of carbon – 14 to be 5715 years?
1
Expert's answer
2020-05-27T13:35:57-0400

1/16 means 4 half-lives:

1/2 * 1/2 * 1/2 * 1/2 = 1/16

4 * 5715 = 22 860 years